• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 19:26:14 on October 30,1998

    Entry number 4200

    keyword=target NMR

    18:27 UP: -14.47mV (fitted)
    DN: -14.49mV (fitted)
    --->39.08% EPR
    --->37.64% H20

    --------------------------------------------
    18:23 target in NMR position (20022)
    18:27 NMR/RTD screen shot
    18:30 lasers turned off. target to 39238
    18:31 began rot to |-
    18:32 transvers laser accidently activated for roughly 2 seconds
    when laser 7 was brought out of Standby.
    18:36 Rotation complete. turned on laser #4
    18:38 laser #7 on
    18:40 laser #8 on
    18:42 All lasers turned off as Bogdan asks for BeO raster run.
    18:53 target back in postion (39244). Laser 4 back on.
    18:55 laser 7 back on.
    18:57 laser 8 back on.